OpenStudy (anonymous):

Brad's age is five more than four times John's age. The sum of their ages is 25. How old is John?

OpenStudy (anonymous):

Each sentence represents an equation. if B is Brad's age, and J is John's age, what is the equation represented by the statement "Brad's age is 5 more than 4 times John's age"?

OpenStudy (anonymous):

so he is 5

OpenStudy (anonymous):

No. B = 5 + 4*J (Brad's age) (is) (5 more) (4 times Johns age)

OpenStudy (anonymous):

Try this one. "The sum of their ages is 25"

OpenStudy (anonymous):

21 -5 = 16 / 4 = 4 so John is 4 and Brad is 21 - 21+4 = 25